There are no relationship between this 3 events, the probability to give birth to a girl can not be deduce from previous birth given by a women.
So the probability that the third child will be a girl is the same for this woman as for any other woman, no matter what happened before.
It's like rolling a dice, if one gets an ace and an ace, it does not change the chance to get another ace for the third roll, or a 2, or a 3,....
All the roll results are independent if the dice is not loaded
The random variable z is the standard normal score.find the p -aza equals 0.6600?
Unfortunately, limitations of the browser used by Answers.com means that we cannot see most symbols. It is therefore difficult to be sure of giving a proper answer to your question.
Assuming the question was to find a such that Prob(-a < z < a) = 0.6600, the answer is a = 0.954
If the question was something else, please resubmit your question spelling out the symbols as "plus", "minus", "times", "equals".
How many different ways to arrange the word critics?
The letters of the word critics can be permuted in 7!/2! = 2520 ways.
How can doctors and genetic counselors calculate the probability of a child inheriting a disease?
the punnett square
As long as the number of gum balls is "large", the probabilities will change very litle with only one candy removed.
If the probability of flipping a coin is one half than why don't you get heads half the time?
The probability is the split you expect in an infinite number of trials.
Every flip has a probability of 50/50 regardless of what came before it.
Even if the previous 972 flips all came up heads, there's no such thing as
"It's due for a tails", because a coin doesn't remember anything.
So there are two possible easy answers to your question:
1). A real coin, with different engraving on the 2 sides, and maybe some dirt or grease on it too,
is not perfectly balanced, and its probability is not exactly 50/50. In other words, every real coin
is slightly loaded.
2). You didn't flip it enough times for the split to be statistically significant, that is, for any
probabilistic tendency to emerge. Try 100 flips some rainy Saturday afternoon and see
what you get.
What is the probability that something will never happen?
It depends on what the "something" is.
The probability that the roll of a regular die will never exceed 6 is 1.
The probability that the roll of a regular die will never exceed 3 is 0.5.
The probability that the roll of a regular die will never exceed 0 is 0.
What the probability of an odd number?
The answer depends on what set of integers is under consideration.
The answer depends on what set of integers is under consideration.
The answer depends on what set of integers is under consideration.
The answer depends on what set of integers is under consideration.
How is science related to basketball?
Like when i be in math and English class i have to learn different kinds of problems and different kinds OS sentences,verbs and etc.School is so exciting and fun cause you have friends and people there for you and people who teaches you something and make sure you eat and make sure you learn something so you can go home and tell your parents how your day went.Like when your teachers send homework home don't't give it to you so you can not do it , let it just be in your bookbag or make up a fib and say your dog ate they suspect you to do it and turn it in the next day not next week,next month or even next year that will just be a flat F.SO THAT IS MY STORY ABOUT HOW SCIENCE RELATE TO BASKETBALL.
What is the probability of not rolling a 8 on dice?
0n a 6 sided dice. exactly 1. on two 6 sided dice, about 31/36
What is the theoretical probability of spinning a 3 and picking an a?
There is insufficient information in the question to properly answer it. You did not specify how many places and how many 3's there are on the spinner, and you did not specify the set from which you could pick an "a", both of which are necessary pieces of information to calculate probability. Please restate the question.
Relationship between Exponential and Poisson Distributions?
Poisson distribution shows the probability of a given number of events occurring in a fixed interval of time. Example; if average of 5 cars are passing through in 1 minute. probability of 4 cars passing can be calculated by using Poisson distribution.
Exponential distribution shows the probability of waiting times between occurrences of events.
If we use the same example; probability of a car coming in next 40 seconds can be calculated by using exponential distribution.
-Poisson : probability of x times occurrence
-Exponential : probability of waiting times between events.
Does standard deviation answer have to be between 3?
No, since "between 3" doesn't mean anything. Do you mean between -3 and +3?
What is the probability of rolling more than a six?
The probability of rolling more than a six with one roll of a cube whose sides are marked from 1 to 6 is zero.
What does or mean in probibility?
"or" is used in the context of sets [of events] rather than probability (and certainly not probibility!),
An event described as A or B means either event A or event B or both events.
"or" is used in the context of sets [of events] rather than probability (and certainly not probibility!),
An event described as A or B means either event A or event B or both events.
"or" is used in the context of sets [of events] rather than probability (and certainly not probibility!),
An event described as A or B means either event A or event B or both events.
"or" is used in the context of sets [of events] rather than probability (and certainly not probibility!),
An event described as A or B means either event A or event B or both events.
How many arrangements can be made from 6 pictures on a wall?
6! = 720 would be the number of different arrangements you can make from 6
different pictures.
If you pull 35 marbles without replacement, the answer is 1: the event is a certainty.
If you pull only one marble, at random, the probability is 16/50 = 8/25.
What does it mean for two events to be dependent?
An even is dependent of the other when one has to occur before the other can do so. Else, one can also occur with the other.
E.g. Christmas (or the winter holidays if you will) has to happen/so that/you go buy Christmas presents.
Hunger usually "depends" on the lack of food in the stomach.
---
In math, events are dependent when one event affects the probability of the other.
E.g. There is 1 red, 1 green and 1 blue balls in a container. You pick a ball out of 3. There is a probability of 1/3 = 33% of being a certain colour. Now there are only two balls in the container. The next one you pick has a probability of 1/2=50% of being a certain colour,
What world events were occurring in the year 1914?
all i know is that 2 yrs before that the titanic sank
What are the types of probability?
They are generally agreed to be theoretical and experimental probabilities.
Probability is probability. The concept may be applied to any causal event which has more than one potential outcome.